### v1

# ppt: A Process-based Preference Learning Framework for Self Improving Table Question Answering Models

Abstract
Improving large language models (LLMs) with self-generated data has demonstrated success in tasks such as mathematical reasoning and code generation. Yet, no exploration has been made on table question answering (TQA), where a system answers questions based on tabular data. Addressing this gap is crucial for TQA, as effective self-improvement can boost performance without requiring costly or manually annotated data. In this work, we propose ppt, a Process-based Preference learning framework for TQA. It decomposes reasoning chains into discrete states, assigns scores to each state, and samples contrastive steps for preference learning. Experimental results show that ppt effectively improves TQA models by up to 5% on in-domain datasets and 2.4% on out-of-domain datasets, with only 8,000 preference pairs. Furthermore, the resulting models achieve competitive results compared to more complex and larger state-of-the-art TQA systems, while being five times more efficient during inference.
1 Introduction
Table Question Answering (TQA) aims to answer questions based on tabular data. Current approaches developing TQA models Wu and Feng (2024); Zhang et al. (2025b) fine-tune pre-trained large language models (LLMs) on labeled answers augmented with Chain-of-Thought (CoT) reasoning Wei et al. (2023).
Figure 1: Given a TQA problem, a TQA model generates reasoning states, composed of cumulative reasoning steps. A value function is used to score each state. We roll out parent states to obtain pair-wise steps and select those with value differences larger than a specified threshold. After data collection, we perform step-wise preference learning for improving the TQA model.
Recent studies suggest that LLM performance can be improved by fine-tuning with self-generated data on many tasks, such as mathematical reasoning Singh et al. (2024) and code generation Zelikman et al. (2024). Yet, no exploration of self-improvement has been conducted in TQA. Addressing this gap is crucial, as an effective improvement method further enhances an existing model without costly or manually annotated data.
To self-improve TQA models, two challenges arise: (1) devising an effective learning strategy and (2) automatically obtaining high-quality data for learning. Drawing inspiration from contrastive process-based learning Lai et al. (2024); Zhang et al. (2025a), where pairs of intermediate reasoning steps provide contrastive signals for model learning, we propose ppt, a novel Process-based Preference learning framework for TQA.
As shown in Figure 1, our framework contains three stages: state generation, state value estimation and pair-wise step sampling. During the first stage, we sample multiple reasoning chains from a TQA model and formulate them into reasoning states, composed of cumulative reasoning steps, During state value estimation, a value function gives continuous scores based on the potential of a state for deriving a final correct answer. During pair-wise steps sampling, we roll out parent states to obtain paired child states and select state pairs with value differences larger than a specified threshold. After obtaining preference data, we perform process-based preference learning using direct preference optimization (DPO) Rafailov et al. (2024) to self-improve the TQA model.
Experiments show that ppt effectively improves TQA models by up to 5% on in-domain datasets and 2.4% on out-of-domain datasets with only 8k preference data. It achieves similar performances compared to methods requiring additional LLMs as judges, yet being ten times more efficient. The enhanced model outperforms existing fine-tuned TQA models. On three datasets, it even achieves similar performances compared with much larger and complex frameworks, yet boasts five times more efficiency during inference.
2 Related Work
We discuss current approaches in addressing TQA and methods for self-improvements.
Table Question Answering.
Current TQA systems can be categorized into two groups: fine-tuned models Zhang et al. (2023a); Wu and Feng (2024) and training-free frameworks Zhou et al. (2025); Nahid and Rafiei (2024). The former collects existing TQA datasets to train an end-to-end system that either directly outputs answers Zhang et al. (2023a) or makes use of an additional reasoning chain Wu and Feng (2024). small-size models ($\leq 8B$) are utilized out of efficiency considerations. Training-free frameworks in TQA are dependent on larger models, e.g., GPT-3.5, Qwen-72B Qwen et al. (2025) and involve more complex designs, such as tools Nahid and Rafiei (2024) and agentic collaborations Zhou et al. (2025). We focus on the fine-tuning approach given its inference efficiency.
Self-Improving LLMs.
Self-improving focuses on using training data generated by the model itself to further obtain performance gains Singh et al. (2024); Zelikman et al. (2024); Xiong et al. (2024); He et al. (2024). Common methods applied for fine-tuning include rejected sampling fine-tuning (RST) Yuan et al. (2023) and DPO Tu et al. (2025); Xu et al. (2024); Zhang et al. (2025a). To obtain step-DPO data for preference learning, previous work either uses closed-source models for locating errors Lai et al. (2024) or existing process supervision datasets Lightman et al. (2023) to train a step reward model. Zhang et al. (2025a). In contrast, our framework does not involve LLM judges, nor does it depend on manually annotated process supervision data. To the best of our knowledge, we are the first to explore process-based supervision for self-improvement in TQA.
3 Process-based Preference Learning for TQA
Task Setup.
Given a table $t$, a question $q$, a fine-tuned TQA model $M_{ft}$ should provide a correct answer $a$. To enhance systems’ performances, a reasoning chain $r$ is generated along with $a$. $r$ consists of $m$ steps: $\{k_{1},k_{2},...,k_{m}\}$. A problem ($t,q$) and an instruction ($u$) form an initial state $s_{0}$. Each new state $s_{i}$ consists of previous state $s_{i-1}$ and a step $k_{i}$ generated at the timestep $i$: $s_{i}=(s_{i-1},k_{i})$ and $s_{i-1}=\{s_{0},k_{1},k_{2},....k_{i-1}\}$.
State Generation.
Given a TQA problem, we first sample $m$ reasoning chains from $M_{ft}$, resulting in $\{r\}_{i=1}^{m}$. Prompts can be found in Appendix A.1. Each reasoning chain consists of multiple steps. We accumulate steps to construct states: $s_{i}=\{s_{0},k_{1},...k_{i}\}$. If all reasoning traces lead to correct answers, we abandon the problem as it might be too easy and only lead to high value states.
State Value Estimation.
A state value function $V$ takes in a state and returns its value. We approximate a state’s value by Monte Carlo (MC) sampling: $V(s_{i})$ is calculated as the probability of $s_{i}$ leading to the correct answer. More specifically, $M_{ft}$ takes in $s_{i}$ and completes the current reasoning chain until reaching an answer. This is repeated $n$ times, resulting in $\{(s_{i+1,j},...,s_{z_{j},j},a_{j})\}_{j=1}^{n}$, where $a_{j}$ and $z_{j}$ are the predicted answer and the total number of steps for the $j$-th finalized solution, respectively. $V(s_{i})$ can be represented as:
$V(s_{i})=\frac{\sum_{j=1}^{n}\mathbb{I}(a_{j}=a)}{n}$ (1)
The continuous value allows more flexible selection of pair-wise steps described in the following paragraph as we can control the distance between the selected states.
Pair-wise Step Sampling
After obtaining values for each state, we filter out intermediate states $s_{i}$ where $V(s_{i})=0$ and also their child state $\{s_{i+1},...s_{z}\}$. This is exemplified by the red nodes in Figure 2. We assume a state of value 0 to contain erroneous steps. Rolling out from it is likely to create bad quality child states. For each remaining state $s_{i}$, we use the completion traces sampled when calculating $V(s_{i-1})$: $\{(s_{i,j},...,s_{z_{j},j},a_{j})\}_{j=1}^{n}$ as roll outs, and calculate state values for each sampled $\{s_{i,j}\}_{j=1}^{n}$. This results in a set of pair-wise states: $(s_{i},\{s_{i,j}\}_{j=1}^{n})$ that can be used to construct step-wise preference dataset $D_{sdpo}$. As $V(s_{i})$ is a continuous value, a good state $s_{w}$ and a bad state $s_{l}$ pair is selected if $V(s_{w})-V(s_{l})\geq\tau$, where $\tau$ is a hyper-parameter. The preference data for step DPO can be represented as $D_{sdpo}=\{(s_{i-1},k_{i}^{w},k_{i}^{l})^{d}\}_{d=1}^{|D_{sdpo}|}$.
Preference Training.
After gathering the process-based preference dataset, we fine-tune $M_{ft}$ using pairs of good and bad steps given previous steps. The loss function is:
$\mathcal{L}_{\text{SDPO}}=-\mathbb{E}_{(s_{i-1},k_{i}^{w},k_{i}^{l})\sim\mathcal{D}_{sdpo}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(k_{i}^{w}|s_{i-1})}{\pi_{ref}(k_{i}^{w}|s_{i-1})}\right)\right.\\ \left.-\beta\log\frac{\pi_{\theta}(k_{i}^{l}|s_{i-1})}{\pi_{ref}(k_{i}^{l}|s_{i-1})}\right]\end{aligned}$
4 Experiments
TQA Models.
Note that existing fine-tuned TQA systems do not feature clear step separation in their reasoning chains. Separating steps by line breaks as in math Luo et al. (2024) is problematic since a line break also indicates a table row in TQA. Therefore, we fine-tune TQA models ourselves that can output clearly separated steps. Details are presented in Appendix A.3. We term the fine-tuned model $M_{ft}$.
Baselines.
We consider baselines for self-improvement strategies: RFT Yuan et al. (2023) trains a model with correct reasoning traces using supervised fine-tuning. FDPO Xu et al. (2024) trains a model with pair-wise correct and incorrect full reasoning chains using DPO. Baselines for value functions include: MC with binary labels (mc-b), it returns binary state values based on if states derive final correct answers. Mixed estimation (mix) Zhang et al. (2025c) scores $s_{i}$ 1 if both mc-b and an external LLM judge $M_{j}$ output 1. If both judges return 0, $s_{i}$ is 0. States receiving different scores by the judges are not considered for building the preference dataset. This accounts for approximately 36% of data (Figure 5). Baselines for TQA models include both end-to-end Zhang et al. (2023a); Wu and Feng (2024) and training-free frameworks Zhou et al. (2025); Nahid and Rafiei (2024). Details of the base methods and models can be found in Appendix A.4.
Figure 2: Process-based preference data collection. We estimate a state value by the probability of a state leading to a correct answer. In the first example, $V(s_{i})=\frac{2}{3}$. After obtaining state values, we do not consider intermediate states that have a value of 0 ($s_{21}$), together with their child states ($s_{31}$). We sample pair-wise states for each remaining state, e.g., $s^{\prime}_{22}$ is sampled by rolling out $s_{12}$ and is regarded as a pair state for $s_{22}$.
Datasets.
We train $M_{ft}$ using the training sets of three common TQA datasets: WTQ Pasupat and Liang (2015), TabFact Chen et al. (2020), and HiTab Cheng et al. (2022). To obtain preference data, we sample from their validation sets. We use the test sets of these three datasets as in-domain evaluation data, and incorporate three out-of-domain datasets: WikiSQL (Zhong et al., 2017), SCITAB Lu et al. (2023), and CRT Zhang et al. (2023b) to test models’ generalisability. Details about the datasets are presented in Appendix A.2.
Experimental Settings.
We choose Qwen-2.5-7B Qwen et al. (2025) and LlaMA-3.1-8B Grattafiori et al. (2024) as base models. During preference learning, we fix the fine-tuning dataset size to 8k for Qwen-2.5-7B and 6.7k for LlaMA-3.1-8B, as different baselines result in different sample sizes (statistics are shown in Appendix A.5). For fair comparison, we use the smallest sample size collected as the fine-tuning data size. We make sure all sampled data correspond to the same set of TQA problems to rule out instance-level effects on fine-tuning. Hyper-parameters are shown in Appendix A.6. We use Qwen-2.5-72B as $M_{j}$. The Prompt for $M_{j}$ is shown in Appendix A.1. The number of reasoning chains $m$ is set to 4 and the roll-out number $n$ is set to 8. The threshold $\tau$ is set to 0.9 and temperature is set to 0.7 and 0 during dataset construction and inference respectively. We use Exact Match as the evaluation metric. All experiments are conducted using 4 A100 GPUs. Training is performed with LlaMA-Factory Zheng et al. (2024) and inference is performed with VLLM Kwon et al. (2023).
5 Results and Discussions
Figure 3 shows the Exact Match of different methods on in-domain datasets, averaged across models. Per model results can be found in Appendix A.7.
Figure 3: Comparing our ppt with baselines using Exact Match (EM). Results are averaged across models. RFT and FDPO stand for rejected sampling fine-tuning and full chain DPO. We experiment with value functions: mc-b and mix. Dashed lines show performances of fine-tuned TQA models $M_{ft}$ before applying self-improvement methods.
ppt effectively improves performance of TQA models.
As Figure 3 shows, applying ppt enhances the performance of $M_{ft}$ by 3.5% on average on in-domain datasets. The gains are more obvious on WTQ (5%) compared to TabFact (2%). This might be attributed to dataset features: TabFact is a binary classification dataset, thus it is easier for models to achieve high performance and harder to be further improved. When evaluating on out-of-domain datasets, we witness an average of 2.2% performance gain after applying our framework ($\uparrow$ 2.2% for WikiSQL, $\uparrow$ 2.4% for SCITAB and $\uparrow$ 2.1% for CRT). The findings show the genealisibility of our method on out-of-domain data.
Process-based preference learning is effective and efficient for self-improving TQA models. Comparing ppt with RFT and FDPO, we find RFT does not improve performance. In contrast, preference learning effectively self-improves models. Moreover, process-based preference learning (ppt) outperforms coarse-grained full chain learning (FDPO) on two out of three datasets. Comparing different value functions, we find ppt with continuous MC estimation to work better than other methods. This is an important finding as it shows that simply using $M_{ft}$ to estimate state values with a high threshold $\tau=0.9$ to control pair qualities can work similarly or better than additionally querying an LLM judge. It significantly boosts the efficiency of the framework: we find using ppt takes 10 times less time than using ppt(mix) when sampling the same number of data. In a manual evaluation of step correctness (Shown in Appendix A.8), we do not find differences between ppt and ppt(mix). This suggests ppt’s effectiveness and efficiency are not compromised for reasoning correctness. We also provide an analysis of threshold impact in Appendix A.9, showing the necessity of picking a relatively high threshold for distinguishing step pairs during preference learning.
Models WTQ TabFact HiTab WikiSQL SCITAB CRT
Protrix 56.2 71.6 - 67.4 45.0 40.2
T-LlaMA 35.0 82.6 64.7 50.5 38.6 26.9
$M_{ft}$ 54.9 82.4 61.4 68.3 54.9 48.5
$M_{si}$ 63.1 84.9 67.6 72.0 56.9 51.4
MACT 70.4 - - - 55.8 57.4
T-SQLify 64.7 80.2 - 76.7 50.9 42.0
Table 1: Exact Match of TQA models. $M_{ft}$ and $M_{si}$ refers to fine-tuned and self-improved TQA models respectively. T-LlaMA refers to TableLlaMA and T-SQLify refers to TabSQLify. Framework results (the last two rows) are obtained using GPT-3.5 as the backbone. SoTA TQA results are obtained from previous work Zhou et al. (2024); Zhang et al. (2023a); Wu and Feng (2024); Nahid and Rafiei (2024).
Self-improved TQA models achieve competitive results compared with more complicated SoTA approaches with five times less inference time. Table 5 shows results for current TQA models. For the sake of space, we only show results from Qwen-2.5. Full results containing LlaMA-3.1 are shown in Appendix A.7. The first four rows of Table 5 compare small-size fine-tuned TQA models, and the last two rows show SoTA training-free frameworks back-boned by GPT-3.5. We find $M_{si}$, enhanced using ppt outperforms existing TQA models. More importantly, it achieves competitive performance compared with larger and more complex frameworks where tools and agentic collaboration are involved. On SCITAB, $M_{si}$ even achieves the best performance. We emphasize the inference efficiency of $M_{si}$: it requires eight times less inference time than MACT and five times less than TabSQLify, yet provides competitive results.
6 Conclusions
In this study, we provided the first self-improvement framework for TQA, using process-based preference learning. Our framework effectively improved the performance of TQA models up to 5% and 2.4% on in-domain datasets and out-of-domain datasets respectively. The resulting models demonstrated competitive performances compared to state-of-the-art TQA systems, which are dependent on large LLMs and tool usages. Yet, it requires five times less inference time.
Limitation
While our method effectively enhances the performance of small-size TQA models, its impact on large TQA models remains unexplored. To the best of our knowledge, current fine-tuned TQA models only focus on small-size LLMs. Future work can explore an efficient training strategy for large fine-tuned TQA models. Secondly, we limit the task in our study to only TQA, while there exist other table-related tasks, such as table summarization. Thirdly, although our framework supports iterative self-learning, the present work only demonstrates the effectiveness of the first iteration, leaving multi-iteration evaluations for future study. As the datasets we used in this study are originally sourced from Wikipedia, scientific paper and statistical reports, we do not observe any potential risks from the datasets.
Appendix A Appendix
A.1 Prompts
Figure 6, 7 and 8 show prompts for generating a full reasoning trace, completing a reasoning trace and LLM judge evaluation for a reasoning trace.
A.2 Datasets
Table 2 shows number of instances and domains for the test data we used.
A.3 Fine-tune TQA Models
Following previous work Wu and Feng (2024); Zhang et al. (2025b), we prompt Deepseek-V3 DeepSeek-AI et al. (2025) for reasoning chains that lead to correct answers to fine-tune our TQA models. Specially, we sample 2.4k, 1.5k and 2.3k data from the training set of WTQ Pasupat and Liang (2015), TabFact Chen et al. (2020) and HiTab Cheng et al. (2022). Then we prompt Deepseek-V3 to generate reasoning chains and answers. We only keep reasoning chains leading to correct answers for fine-tuning. This results in 1612 instances from WTQ, 1425 instances from TabFact, and 1277 from HiTab. In total, we collect 4314 instances.
A.4 Baselines
Baselines for Self-Improvement Methods.
To collect data for rejected sampling fine-tuning (RFT), we select the reasoning chains leading to the correct answer: $D_{rft}=\{(r_{w})^{i}\}_{i=1}^{|D_{rft}|}$. We fine-tune TQA models with the following loss function for self-improvement:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(s_{0},r_{w})\sim\mathcal{D}_{rft}}\left[\log\pi_{\theta}(r_{w}|s_{0})\right]$ (3)
To obtain full reasoning chain DPO data, we selecting reasoning chains leading to correct as preferred traces and those leading to incorrect answers as not preferred: $D_{fdpo}=\{(s_{0},r_{w},r_{l})^{i}\}_{i=1}^{|D_{fdpo}|}$, where $r_{w}$ and $r_{l}$ refer to the preferred and not preferred reasoning traces respectively. The loss function can be represented as:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(u,r_{w},r_{l})\sim\mathcal{D}_{dpo}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(r_{w}|u)}{\pi_{ref}(r_{w}|u)}\right)\right.\\ \left.-\beta\log\frac{\pi_{\theta}(r_{l}|u)}{\pi_{ref}(r_{l}|u)}\right]\end{aligned}$
Baselines for TQA Systems.
TableLlaMA Zhang et al. (2023a) is an end-to-end TQA system fine-tuned on 2.6M existing table-related data. It uses LlaMA-2-7B as a base model and directly outputs answers without a reasoning chain. Protrix Wu and Feng (2024) is fine-tuned with around 4k table-related data with reasoning chain generated from GPT-4. It also uses LlaMA-2-7B as the base model. MACT Zhou et al. (2025) is a training-free TQA framework, leveraging tools and agent collaboration. Specifically, a planning agent outputs step-wise plans to decompose questions, and a coding agent and tools solve each step. TabSQLify Nahid and Rafiei (2024) decomposes tables into relevant sub-tables with SQL query generation and execution. Then, sub-tables and questions are passed to LLMs to obtain final answers.
Datasets #instances Domain
WTQ 4344 Wikipedia
TabFact 12779 Wikipedia
HiTab 1584 statistical reports
WikiSQL 15878 Wikipedia
SCITAB 1224 scientific paper
CRT 728 Wikipedia
Table 2: Test data statistics.
A.5 Sampled Dataset Statistics
Table 3 shows sampling size for each method. We find mc-b results in the most data while RFT the least.
Methods HiTab WTQ TabFact Total
Original 1.6k 2.8k 5k 9.4k
RFT 2.2k/1.6k 2.7k/2.4k 3k/2.7k 8k/6.7k
FDPO 3.6k/2k 4.7k/4k 4.6k/3.6K 12.9k/9.6k
SDPO
+mc-b 25k/20.8k 46k/47.1k 37k/38.4k 109k/106k
+mix 5.4k/3.5k 9.6k/9.7k 5k/4.9k 20k/18.1k
+mc-0.9 6.8k/1.6k 16k/4.4k 11k/2.4k 33.8k/8.4k
Table 3: Sampled dataset sizes for different methods. Results for Qwen-2.5-7B and LlaMA-3.1-8B are separated by “/”. mc-b refers to using Monte Carlo sampling with binary values as the value function. mix stands for using both mc-b and an LLM judge (Qwen-2.5-72B) as the value function. mc-0.9 stands for using Monte Carlo sampling with continuous values and setting the selection threshold as 0.9. RFT refers to rejected sampling fine-tuning. FDPO and SDPO stand for full-trace DPO and step-wise DPO.
A.6 Hyper-parameters
Table 4 shows the hyper-parameters used for model fine-tuning.
Models Method Fine-tuning Learning rate Epoch Batch size LoRA rank DPO $\beta$
Qwen-2.5-7B supervised fine-tuning full-parameter 5e-6 2 128 - -
Qwen-2.5-7B rejected sampling fine-tuning LoRA 1e-5 1 128 64 -
Qwen-2.5-7B full chain DPO LoRA 1e-5 3 128 64 0.1
Qwen-2.5-7B step-wise DPO LoRA 1e-5 3 128 64 0.1
LlaMA-3.1-8B supervised fine-tuning full-parameter 5e-6 2 128 - -
LlaMA-3.1-8B rejected sampling fine-tuning LoRA 1e-5 1 128 32 -
LlaMA-3.1-8B full chain DPO LoRA 1e-5 3 128 32 0.1
LlaMA-3.1-8B step-wise DPO LoRA 1e-5 3 128 32 0.1
Table 4: Hyper-parameters used for model fine-tuning.
A.7 Additional Results
Table 5 shows different models’ performance on the six investigated datasets.
Models WTQ TabFact HiTab WiKiSQL SCITAB CRT In-domain Out-of-domain
Qwen-2.5-7B 28.66 73.77 26.07 47.76 39.46 35.58 42.83 40.93
+ TQA training ($M_{ft}$) 54.93 82.37 61.36 68.26 54.90 48.49 66.22 57.22
       +RFT 56.26 82.44 60.29 68.15 52.29 46.15 66.33 55.53
       +FDPO 61.10 84.65 63.83 72.04 56.94 52.06 69.86 60.34
       ppt(mc) 60.80 82.98 65.72 70.08 52.04 50.24 69.83 57.45
       ppt(mix) 63.86 85.32 64.20 71.44 55.56 52.47 71.63 59.82
       ppt ($\tau=0.9$) 63.10 84.88 67.55 71.97 56.94 51.37 71.84 60.09
LlaMA-3.1-8B 30.64 63.91 26.20 31.87 43.38 32.55 40.25 35.93
+ TQA training ($M_{ft}$) 64.80 83.36 69.63 69.48 51.63 49.04 72.60 52.76
       +RFT 62.06 84.15 67.30 70.73 50.49 48.63 70.84 56.62
       +FDPO 65.22 85.91 68.18 71.34 53.35 50.41 73.10 58.37
       ppt(mc) 65.54 84.01 69.95 70.67 50.25 49.31 73.17 56.74
       ppt(mix) 62.39 84.98 66.79 68.31 52.94 51.24 72.05 57.50
       ppt ($\tau=0.9$) 65.84 85.50 70.33 70.08 54.44 50.27 73.98 58.26
Table 5: Exact Match accuracies of models fine-tuned with different strategies and value functions, generated with greedy decoding. We bold the best results and underline the second best results for each model type.
A.8 Reasoning Chains Analysis
We sample 100 reasoning chains leading to correct answers generated from models using ppt and models using mix as the value function. We manually examine the correctness of the reasoning chains. Among the 100 instances, we exclude 8 instances where either the answers are incorrect or the questions are ambiguous. We find similar accuracies of the reasoning chains generated from the aforementioned methods, with 95.7 and 94.6, respectively. This suggests the two methods do not differ much in terms of leading to correct reasoning chains. Nevertheless, wrong reasoning chains leading to correct answers still exist, possibly due to overly complex table inputs. An error case is shown in Figure 9.
WTQTabFactHiTabWikiSQLSCITABCRT5050505055555555606060606565656570707070757575758080808085858585DatasetsEMThresholds Comparison with Soft EstimationSoft Estimation ($\tau$=0.5)Soft Estimation ($\tau$=0.7)Soft Estimation ($\tau$=0.9)
Figure 4: Thresholds comparisons with different value functions on six TQA datasets.
Figure 5: Confusion matrix between LLM judge evaluation and Monte Carlo estimation with binary labels.
A.9 Threshold Analysis
The threshold $\tau$ decides the state value differences when sampling a pair of (preferred and not preferred) states. We set $\tau$ to 0.9 in our study. We experiment with different values of $\tau$ to investigate its impact on the fine-tuned process-supervised models. The experimental settings are the same as described in Section 4 except that we change the values of $\tau$. Figure 4 shows the performances of models fine-tuned with data sampled using different $\tau$. We observe that there is a tendency for higher thresholds to lead to better performance. However, we do not observe big differences in terms of model performances when setting $\tau$ to 0.7 or 0.9.
A.10 Dataset Licenses
WTQ Pasupat and Liang (2015), HiTab Cheng et al. (2022) and WikiSQL Zhong et al. (2017) are under the license of CC-BY-SA-4.0111https://creativecommons.org/licenses/by-sa/4.0/, BSD-3 Clause [^2] https://opensource.org/license/bsd-3-clause and C-UDA [^3] https://github.com/microsoft/HiTab?tab=License-1-ov-file respectively. TabFact Chen et al. (2020), CRT Zhang et al. (2023b) and SCITAB Lu et al. (2023) are under the MIT [^4] https://opensource.org/license/mit license.
Figure 6: Prompt to generate full reasoning trace given a TQA problem.
Figure 7: Prompt to complete a reasoning trace given a TQA problem and past steps.
Figure 8: LLM judge prompt to generate analysis, decisions and confidence for each step.
Figure 9: Wrong reasoning chain generated by a self-improved model (Qwen-2.5-7B) using ppt. The first wrong step is highlighted with red.